OpenStudy (anonymous):

How do I simplify: 2m/m-9 + 18/9-m?

OpenStudy (anonymous):

HELP! D:

OpenStudy (anonymous):

\[\frac{2m}{m-9} + \frac{18}{9-m}\]

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

then what happens? cos I know I have to do this: 9-m = --9+-m = -(-9 + m) = - (m-9)

OpenStudy (anonymous):

common denominator Multiply -1 at top and bottom of the first fraction

OpenStudy (anonymous):

I multiply -1 because I have a -1 on my -(m-9(?

OpenStudy (anonymous):

\[\frac{-2m}{9-m} +\frac{18}{9-m}\]

OpenStudy (anonymous):

factor 2 from -2m + 18 2(-m + 9)/ (9-m) answer is 2

OpenStudy (anonymous):

what did you do?

OpenStudy (anonymous):

\[\frac{-2m+18}{9-m}\]

OpenStudy (anonymous):

and you factor -2m + 18?

OpenStudy (anonymous):

\[\frac{2(9-m)}{9-m}\]

OpenStudy (anonymous):

yeah factor the two

OpenStudy (anonymous):

ohhh, okay. thank you!